OpenStudy (anonymous):

The measures of one of the acute angles in a right triangle is given. Find the measure of the other acute angle. given :(90-z)

OpenStudy (anonymous):

a+b+c=180 we know that one of the angles is 90, let's say x let's call y the given angle 90+(90-z)+c=180 180-z+c=180 that means that the 3rd angles has to be z c=z

Directrix (directrix):

The acute angles of a right triangle are complementary. If one is 90 - z, then the other has to be z. 90 - z + z = 90.

OpenStudy (anonymous):

thanks @Directrix

Directrix (directrix):

You are welcome.